# Dynamical friction and measurements of the splashback radius in galaxy clusters

### v1

@@ 3 Results
3.1 Residence Time
The splashback radius is traced out by subhaloes that have spent sufficient time in the cluster to reach their first apocenter. In position-radial velocity phase space, then, $R_{\mathrm{sp}}$ should appear as an overdensity of subhaloes with low radial velocities. Moreover, galaxies that are just infalling into the cluster should not trace out the splashback radius. In this section, we investigate how the cluster phase-space evolves with residence time.
Figure 6: The ratio of the splashback radius with and without dynamical friction, after 5 Gyr, for the ratio of minimum stellar mass to halo mass (measured in the middle of the bin). The horizontal grey dotted line marks a ratio of 1 (i.e. no difference). The black dashed line shows where $M_{\mathrm{subhalo}}/M_{\mathrm{200,mean}}\approx 0.01$. Finally, the solid grey line and shaded band show average ratio and error for the regions with multiple data points.
The phase-space distributions for high-stellar mass galaxies ($M_{*}>10^{9}~{}\mathrm{M_{\odot}}$) are shown in Figure 7 after 10 Gyr of evolution, for our lowest-mass clusters (between $10^{13}$ and $10^{13.5}\mathrm{M_{\odot}}$). We examine the phase space after 10 Gyr, rather than 5 Gyr like in the rest of the analysis, to see longer-duration changes in the cluster environment. We separate the subhaloes by residence time, which is defined as the time since the subhalo first crossed inside of $R_{\mathrm{200,mean}}$ in 10 Gyr of simulation. The top three panels show subhaloes with residence times of 1 to 3 Gyr, 4 to 6 Gyr, and 7 to 10 Gyr. The bottom panel shows all galaxies meeting the mass cut, including galaxies that never reach within $R_{200,\mathrm{mean}}$ or galaxies that began the simulation inside $R_{200,\mathrm{mean}}$. We filter out a small number of galaxies with $M_{\mathrm{subhalo}}>0.05M_{200,\mathrm{mean}}$, as they should, in a physically realistic environment, influence the movement of the cluster. On the plot they would appear as a horizontal band throughout a range of radial positions with very low radial velocities (approximately zero).
If instead a minimum stellar mass cut of $10^{7}~{}\mathrm{M}_{\odot}$ is applied, the velocity profiles resemble the expected distribution from IllustrisTNG outputs (see e.g. Dacunha et al., 2022; O’Neil et al., 2024), suggesting our model is reasonably realistic, with respect to other simulated clusters (e.g. the large number of infalling galaxies with negative radial velocity, the dispersal of galaxies at long residence times). For galaxies with residence times between 1 and 3 Gyr, we see the infalling and outgoing galaxies combined, as well as a turnaround in phase space around $\sim 5$ Gyr.
The splashback feature is thought to be found among galaxies with residence times of order one crossing time ($\sim$ 5 Gyr, from O’Neil et al. (2024)), with radial velocities around 0. We begin to see these tracer galaxies in the middle residence time bin, and the number grows with residence times of over 7 Gyr (the top panel). While one might expect the splashback feature to be traced by galaxies with residence times of closer to 5 Gyr, the feature is in fact most visible among galaxies that have spent longer in the cluster (residence time between 7 and 9 Gyr). We attribute this phenomenon to weaker gravitational forces in the least-massive clusters producing less plunging orbits.
Figure 7: 2D histograms of radial velocity and radius plots after 10 Gyr of evolution, with (right column) and without (left) dynamical friction, for the lowest mass haloes ($10^{13}$ $10^{13.5}$ M, totaling 770 haloes) and subhaloes with minimum stellar mass $M_{*}~{}=~{}10^{9}$ M. Subhaloes with $M_{\rm subhalo}~{}>~{}0.05M_{200,\mathrm{mean}}$ are also excluded. The top three panels are separated by residence time in the cluster, which is calculated from the first time at which the galaxy reaches a position within $R_{200,\mathrm{mean}}$. The bottom row shows the histogram for all subhaloes with $M_{*}$ > 109 M and $M_{\mathrm{subhalo}}~{}<~{}0.05~{}M_{200,\mathrm{mean}}$, totaling over 47,000 galaxies. The black dashed line shows the median splashback radius with and without dynamical friction (as measured in Figure 5.)
The majority of galaxies in this halo sample never reach inside of $R_{200,\mathrm{mean}}$ (approximately 88% for both with and without DF.) Of the galaxies that do, the influence of dynamical friction is most apparent among the galaxies with longer residence times. There are few differences between the phase-space distributions for galaxies with $t_{\mathrm{res}}<4$ Gyr. By the longest residence time bin, though, the splashback radius (marked with a vertical dashed line) appears slightly reduced. The distribution of radii at $v_{r}\approx 0$ is wider with dynamical friction enabled as well, matching the wider bootstrapped $R_{\mathrm{sp}}$ distributions shown in Figure 8.
A small number of galaxies with $v_{r}<0$ and $74\leq t_{\mathrm{res}}<$ 7 appear only with dynamical friction enabled. By comparison to the phase-space distribution with a lower mass cut, it is apparent that (essentially) only galaxies with high masses exist in this region of phase space, and only when dynamical friction is acting. We can thus see how DF, for a small number of high mass galaxies, dramatically affects their orbital properties (e.g. as visualized in Fig. 1). However, despite these cases, it is also clear that the phase-space distributions are not, overall, dramatically altered with dynamical friction enabled.
3.2 Validation Tests
The use of these idealized simulations, and the analytical implementation of dynamical friction, allows us to uniquely isolate DF. However, the analytical implementation of DF includes approximations not present in an N-body approach. Therefore, in the following section we investigate the robustness of our results.
3.2.1 Bin Widths
When bootstrapping and fitting the density profiles (see Section 2.4), binning is an important criteria to consider. We use the standard logarithmically spaced bins to ensure that the inner bins are well-sampled. Ideally, the splashback radius distributions should not vary with the number of bins used. Independence from binning indicates that the measurements are not residual effects of the binning choices, and instead reflect real variations in measurements of $R_{\rm sp}$. Too small a number of bins may wash out real variations in $R_{\rm sp}$ measurements, while the use of too many bins produces noisy data due to poorly sampled bins.
Various bin densities are used throughout the literature for splashback radius studies. Different sources cover a different radius range, but for studying $R_{\mathrm{sp}}$, they typically go out to $\sim$ 5-10 $R_{\mathrm{200,mean}}$, ranging from 15 logarithmically spaced bins (Murata et al., 2020; Dacunha et al., 2022), to 32 (O’Neil et al., 2022), 40 (McAlpine et al., 2022), and 42 (linear spacing) (O’Neil et al., 2021).
In Figure 8 we show a violin plot of the splashback radius distribution as a function of number of bins used in the bootstrapping and fitting process, to check the resilience of our results to binning. We apply our fitting procedure as outlined in Section 2.4. In the top panel, including all galaxies with stellar mass above $10^{7}~{}\mathrm{M}_{\odot}$, the distributions are narrow and similar between the different binning options (i.e. left to right), indicating we are very accurately measuring the splashback radius. For the bottom panel, the sample of subhaloes is greatly reduced, producing greater error in the bootstrapping and fitting processes. We selected this subset of cluster and subhalo masses as they involved some of the most extreme density profile variation, and most challenging fits. The broader and more varied distributions reflect the greater uncertainties involved. The cluster subset includes 770 haloes, but of the $3.3~{}\times~{}10^{6}$ subhaloes with stellar masses above $10^{7}~{}\rm{M}_{\odot}$, only approximately 21,000 have $M_{*}>10^{10}~{}\mathrm{M}_{\odot}$. Above 32 bins, the distributions appear relatively stable, however, fits judged by eye appeared to decline in quality above 40 bins. Therefore, we use 32 bins for all other data presented throughout.
3.2.2 Constant Coulomb logarithm simulations
As outlined in Section 2.3.1, the $\Lambda$ term in Chandrasekhar’s analytical DF formula is not particularly well-defined. In the previous sections, we have used the default galpy implementation of $\Lambda$, which is computed as in Equation (3). Here, we seek to investigate how strongly the $\Lambda$ term affects the dynamical friction term and the $R_{\mathrm{sp}}$ measurement.
Boylan-Kolchin et al. (2008) and Jiang et al. (2008) found that merging timescales calculated from the analytical Chandrasekhar formulation are underestimated, relative to results from N-body and hydrodynamic numerical simulations, with Jiang et al. (2008) particularly citing the $\Lambda$ term as a source of uncertainty. Thus, we expect our results, if anything, generally overestimate the impact of DF.
In general, some DF implementations instead utilize an unchanging $\Lambda$ related to the scale size of the halo, $R_{\rm halo}$. However such strategies less accurately match results from N-body simulations (Just & Peñarrubia, 2005); specifically, the use of a constant $\Lambda$ produces overly rapid inspirals and unphysical circularization of orbits Hashimoto et al. (2003). Instead, Petts et al. (2015) use a maximum impact parameter of the subhalo’s orbital radius, which forms the basis for the galpy implementation. Petts et al. (2015) note that this method can overly amplify the influence of dynamical friction.
Given these uncertainties, we examine the dependence of our results on $\Lambda$. We do not examine all possible variations of $\Lambda$ implementation, instead focusing on the use of a constant value as studied in Hashimoto et al. (2003); Just & Peñarrubia (2005). They found that the use of a constant $\Lambda$ caused subhaloes to inspiral too quickly, which we might expect to reduce the splashback radius.
The starting distributions of $\mathrm{ln}(\Lambda)$, as calculated in Equation (3), are shown in Figure 9, for subhaloes beginning within $R_{200,\mathrm{mean}}$ at $z=0$. We select only these subhaloes because they are the ones most strongly influenced by dynamical friction (see e.g. Figure 7). We find that ln($\Lambda$) is normally distributed around ln$(\Lambda)\approx 4-5$. The trend towards greater $\Lambda$ among higher-mass clusters may be unintuitive; since we find DF to be less influential in those clusters, we might expect $\mathrm{ln}(\Lambda)$ to be smaller. However, the larger $\Lambda$ distributions among high mass cluster are reflective of the larger scale radii of the more massive clusters, i.e. $R_{200,\mathrm{mean}}$ is larger in more massive clusters. (Note again, we found that $r_{\rm hm}>GM/v^{2}$ for most galaxies in the simulation, so $r_{\rm hm}$ is the overwhelmingly the value used to calculate $\Lambda$, and it does not change greatly with halo mass.) The gravitational forces are simply far stronger in higher-mass clusters, reducing the relative importance of dynamical friction.
To test the impact of $\Lambda$, we implement a set of constant-$\Lambda$ simulations. We evolve all haloes for 10 Gyr as before, but with fixed $\mathrm{ln}(\Lambda)$ values of 2, 5, and 10 for each subhalo. Figure 10 shows how the splashback radius, measured again after 5 Gyr, evolves with $\Lambda$ for selected halo and subhalo masses. As expected, as $\mathrm{ln}(\Lambda)$ increases, the splashback radii decrease, especially for low-mass haloes. More massive haloes see little impact from adjustments of $\Lambda$, even when they exceed the median starting value by more than a factor of two. We attribute this to the relative weakness of DF compared to gravitational forces, such that significantly increasing DF has very little effect on the overall splashback distribution. Interestingly, even with a $\mathrm{ln}(\Lambda)=2$, which is smaller than the starting median value, the splashback radius is slightly reduced (though not significantly). That finding is consistent with previous work stating that using a constant $\Lambda$ contributes to overly rapid inspiral and enhanced dynamical friction (Hashimoto et al., 2003; Just & Peñarrubia, 2005). Nonetheless, our simple examination into the complexities and uncertainties in defining $\Lambda$, suggest that variations in $\Lambda$ cannot easily account for the splashback radius discrepancies seen in high-mass clusters.
Figure 10: Median splashback radius distributions for constant $\Lambda$ systems, after 5 Gyr of evolution. Low and high mass haloes are included (purple and orange lines, respectively, as marked), with various stellar mass cuts (above $10^{7}$ M in solid lines and $10^{9}$ M dashed). Systems with no dynamical friction are shown at $\mathrm{ln}(\Lambda)$ = 0. The lightened lines show the position of the median splashback radius from the default variable implementation of dynamical friction (as used in Figures 2 to 9.) The shaded bands show the 16-84 percentile range for all data.
@@ 4 Conclusions
Values of the splashback radius, when measured through galaxy number density, have been found to be smaller than splashback radius measurements made from dark matter profiles (on $\gtrsim 10$% scales). This trend persists between observations and simulations. Dynamical friction has been proposed as a force that would uniquely impact the orbits of galaxies, without affecting the dominant, but diffuse, background components of the halo.
@@
  • The Coulomb logarithm term, ln$(\Lambda)$, is a source of uncertainty in the Chandrasekhar dynamical friction implementation. To test the resilience of our results, instead of adjusting $\gamma$, we artificially adjusted, and fixed, $\Lambda$. We find that increasing this logarithm reduces the splashback radius, because it strengthens the dynamical friction force. However, we observe significant changes in the splashback radius only for fixed ln$(\Lambda)$ values of 10 in the lowest mass clusters ($10^{13}<M_{\mathrm{200,mean}}/\mathrm{M_{\odot}}<10^{13.5}$), which is over twice the median starting value of ln$(\Lambda)$ for those clusters (see Figs. 9 and 10). Thus, our results appear consistent even with some variation in ln$(\Lambda)$. High-mass clusters are even less responsive to adjustments in ln$(\Lambda)$.
@@
The semi-idealized simulations presented here exclude other processes that affect the orbits of subhaloes inside the cluster and alter the masses of these subhaloes. For example, processes such as tidal stripping, if they were included, are expected to reduce the impact of dynamical friction as they reduce the masses of subhaloes during infall (van den Bosch et al., 2016, e.g.). Our use of Chandrasekhar’s analytical formula may also overestimate the strength of the dynamical friction (Boylan-Kolchin et al., 2008; Jiang et al., 2008). Each of these factors would further reduce the contribution of DF to splashback radius discrepancies measured in observation or from cosmological simulations.
This investigation targets isolated clusters at $z=0$. It is well-known that DF plays a significant role in halo mergers, which we do not study here. Measurements of halo splashback radii at higher redshift, with higher concentrations (e.g. Adhikari et al., 2016) or among less relaxed clusters, may be more strongly affected by dynamical friction.
@@ Acknowledgements
The authors thank Jo Bovy for their help with galpy, and for their rapid responses to questions raised on the galpy GitHub page. The authors also would like to thank Frank van den Bosch for productive discussions during the early stages of the project. TMO acknowledges support through the Beckman Scholars Program by the Arnold and Mabel Beckman Foundation. TMO also thanks the efforts of her Wellesley College advisor James Battat for his support throughout the research process. JB acknowledges support from NSF grant AST-2153201.

### v2

@@ 4.5 Splashback radius measurements
Despite the large number of recent studies on the splashback radius, the question of why galaxies trace smaller $R_{\mathrm{sp}}$ values than the dark matter has remained unclear. Many studies have attributed this lag to DF, but few have investigated the process itself. Our simulations, idealized though they are, provide a unique window into the behavior of dynamical friction in the cluster environment. The uncertainties inherent to our methodology will enhance the impact of dynamical friction (e.g. the dynamical friction formula, the lack of tidal stripping), suggesting that dynamical friction is even weaker than suggested by our findings here. We thus find that dynamical friction is an insufficient answer to these discrepancies. Future studies of the splashback radius are challenged to develop a more comprehensive explanation for the splashback radius discrepancies (whether it be due to selection effects or limitations in the splashback radius profile fitting, as outlined by O’Neil et al., 2022; Diemer, 2024, respectively).
@@ 5 Conclusions
Values of the splashback radius, when measured through galaxy number density, have been found to be smaller than splashback radius measurements made from dark matter profiles (on $\gtrsim 10$% scales). This trend persists between observations and simulations. Dynamical friction has been proposed as a force that would uniquely impact the orbits of galaxies, without affecting the dominant, but diffuse, background components of the halo. In contrast to past work, we argue that dynamical friction does not significantly reduce the splashback radius, aside from exceptional circumstances.
@@
The semi-idealized simulations presented here exclude other processes that affect the orbits of subhaloes inside the cluster and alter the masses of these subhaloes. For example, processes such as tidal stripping, if they were included, are expected to weaken the impact of dynamical friction as they reduce the masses of subhaloes during infall (van den Bosch et al., 2016, e.g.). Our use of Chandrasekhar’s analytical formula may also overestimate the strength of the dynamical friction (Boylan-Kolchin et al., 2008; Jiang et al., 2008). Each of these factors would further reduce the contribution of DF to splashback radius discrepancies measured in observation or from cosmological simulations.
This investigation targets isolated clusters at $z=0$. It is well-known that DF plays a significant role in halo mergers, which we do not study here. Measurements of halo splashback radii at higher redshift, with higher concentrations (e.g. Adhikari et al., 2016) or among less relaxed clusters, may be more strongly affected by dynamical friction. Future studies could investigate these parameters, as well as increase the complexity of modeling.
@@ Acknowledgements
We thank the anonymous reviewer for their insightful comments, which contributed greatly to the clarity of the paper. The authors also thank Jo Bovy for their help with galpy, and for their rapid responses to questions raised on the galpy GitHub page. The authors also would like to thank Frank van den Bosch for productive discussions during the early stages of the project. TMO gratefully acknowledges support through the Beckman Scholars Program by the Arnold and Mabel Beckman Foundation. TMO also thanks the efforts of her Wellesley College advisor James Battat for his support throughout the research process. JB acknowledges support from NSF grant AST-2153201.
@@
Appendix A Fitting procedure
To enable analysis of stacked clusters, we perform bootstrapping and fitting of the galaxy number density profiles as described in Section 2.4, based on the Diemer & Kravtsov (2014) profile. In this appendix, we explore two elements of fitting (bin widths, as well as the fit parameters) to ensure the rigor of our fitting methods.
A.1 Bin Widths
When bootstrapping and fitting the density profiles (see Section 2.4), binning is an important criteria to consider. We use the standard logarithmically spaced bins to ensure that the inner bins are well-sampled. Too small a number of bins may wash out real variations in $R_{\rm sp}$ measurements, while the use of too many bins produces noisy data due to poorly sampled bins. Ideally, the splashback radius distributions should not vary with the number of bins used. Independence from binning indicates that the measurements are not residual effects of the binning choices, and instead reflect real variations in measurements of $R_{\rm sp}$.
Various bin densities are used throughout the literature for splashback radius studies. Different sources cover a different radius range, but for studying $R_{\mathrm{sp}}$, they typically reach out to $\sim$ 5-10 $R_{\mathrm{200,mean}}$, ranging from 15 logarithmically spaced bins (Murata et al., 2020; Dacunha et al., 2022), to 32 (O’Neil et al., 2022), 40 (McAlpine et al., 2022), and 42 (linear spacing) (O’Neil et al., 2021).
In Figure 8 we show a violinplot of the splashback radius distribution as a function of number of bins used in the bootstrapping process, to check the resilience of our results to binning. In the top panel, including all galaxies with stellar mass above $10^{7}~\mathrm{M}_{\odot}$, the distributions are narrow and similar between the different binning options (i.e. left to right), indicating we are accurately measuring the splashback radius. For the bottom panel, the sample of subhaloes is greatly reduced, producing greater error in the bootstrapping and fitting processes. The cluster subset includes 770 haloes, but of the $3.3~\times~10^{6}$ subhaloes with stellar masses above $10^{7}~\rm{M}_{\odot}$, only approximately 21,000 have $M_{*}>10^{10}~\mathrm{M}_{\odot}$. We selected this subset of cluster and subhalo masses as they involved some of the most extreme density profile variation. The broader and more varied distributions reflect the greater uncertainties involved. Above 32 bins, the distributions appear relatively stable; however, in order to reduce noise in our data we use 32 bins for all data presented throughout.
A.2 Fitting procedure
As described in Section 2.4, we fit the profile of Diemer & Kravtsov (2014) to our data, following, e.g., O’Neil et al. (2022). This fit involves nine parameters, as well as six demarcations between the inner, transition, and outer fit regions, which are not fit by the fitting implementation.
Figure 9: Median $\chi^{2}$ fit values for the lowest mass clusters (top) and highest mass clusters (bottom). For each cluster mass, fits are performed for each stellar mass cut (columns) and with and without dynamical friction (bottom and top rows respectively). This figure shows the parameter space of $L_{t}$ and $H_{t}$. White dots mark the parameters that minimize the median $\chi^{2}$ for each cluster and galaxy mass subsample.
Diemer (2024) identified issues with the Diemer & Kravtsov (2014) profile, including degeneracies among the parameters. However, we continue to use the Diemer & Kravtsov (2014) profile for consistency with earlier studies of the splashback radius. A further investigation of the fitting method is beyond the scope of this paper. Additionally, because of our idealized methodology, the splashback radius values evolve outwards to larger values than real splashback radii, which the profile was not designed for. Because we are interested in comparing the evolution of $R_{\mathrm{sp}}$ with and without DF, this limitation does not affect our analysis, but we must account for it during fitting. Thus, in this section, we present assessment of our fits.
We examine the parameter space of the inner (lower value $L_{i}$ and higher value $H_{i}$ such that the inner region is fit $L_{i}<r/R_{\mathrm{200,mean}}<H_{i}$), transition ($L_{t}$, $H_{t}$), and outer ($L_{o}$, $H_{o}$) fit regions, running fits on each selected combination for all cluster masses and performing chi-squared tests on each fit. We primarily focus our analysis on the transition region boundaries ($L_{t}$ and $H_{t}$) because those are the areas where we expect the splashback radius to have moved (compared to “physical” clusters for which the fit was derived).
Figure 9 shows the results of the chi-squared tests for our lowest and highest cluster samples, split by galaxy mass and the presence of DF to be consistent with the rest of our analysis. We present only the transition region parameters, $L_{t}$ and $H_{t}$ spaced by 0.2. The highest mass clusters ($10^{14.5}$ M) are significantly more dependent on the fitting boundaries than other cluster masses (of which only one cluster mass is shown for comparison). For the high mass clusters, using $L_{t}=1.2$ is produces significantly better fits.
Not every subsample indicates the same “best-fitting" boundaries. Because the highest-mass clusters display the largest variation in $\chi^{2}$ due to the region boundaries, we choose to use $L_{t}=1.2$ and $H_{t}=2.4$ for consistency. Figure 10 shows two example fits for a single bootstrap in the highest mass cluster sample, demonstrating how increasing $H_{t}$ improves the fit quality and captures a stronger dip in the differential density profile around the splashbak radius. We also use $H_{i}=1.2$ (as opposed to 1.0, as used by O’Neil et al., 2022) so that the inner and transition regions meet. Based on these tests, we thus fit all of the data presented in the main text with an inner region from $0.2<r/R_{\mathrm{200,mean}}<1.2$, a transition region from $1.2<r/R_{\mathrm{200,mean}}<2.4$, and an outer region $2.0<r/R_{\mathrm{200,mean}}<10.0$.
Figure 10: Fits (in light and dark blue) of data points (in green) for one example bootstrap from the highest mass ($10^{14.5}\rm M_{\odot}$) cluster sample. The errors on the bootstrapped data are derived from sub-bootstrapping as per the methodology of Appendix B1 of O’Neil et al. (2022).
Appendix B Coulomb logarithm
As outlined in Section 2.3.1, the $\Lambda$ term in Chandrasekhar’s analytical DF formula is not particularly well-defined. In the main body of text, we have used the default galpy implementation of $\Lambda$, which is computed as in Equation (3). Here, we investigate how strongly the $\Lambda$ term affects the dynamical friction term and the $R_{\mathrm{sp}}$ measurement.
Boylan-Kolchin et al. (2008) and Jiang et al. (2008) found that merging timescales calculated from the analytical Chandrasekhar formulation are underestimated, relative to results from N-body and hydrodynamic numerical simulations, with Jiang et al. (2008) particularly citing the $\Lambda$ term as a source of uncertainty. Thus, we expect our results to, if anything, generally overestimate the impact of DF.
Some DF implementations instead utilize an unchanging $\Lambda$ related to the scale size of the halo, $R_{\rm halo}$. However such strategies less accurately match results from N-body simulations (Just & Peñarrubia, 2005); specifically, the use of a constant $\Lambda$ produces overly rapid inspirals and unphysical circularization of orbits (Hashimoto et al., 2003). Instead, Petts et al. (2015) use a maximum impact parameter of the subhalo’s orbital radius, which forms the basis for the galpy implementation. Petts et al. (2015) note that this method can overly amplify the influence of dynamical friction.
Given these uncertainties, we examine the dependence of our results on $\Lambda$. We do not examine all possible variations of $\Lambda$ implementation, instead focusing on the use of a constant $\Lambda$ value. Past work has found that the use of a constant $\Lambda$ caused subhaloes to inspiral too quickly (Hashimoto et al., 2003; Just & Peñarrubia, 2005), which we might expect to reduce the splashback radius.
[…]
Figure 12: Median splashback radius distributions for constant $\Lambda$ systems, after 5 Gyr of evolution. Low and high mass haloes are included (purple and orange lines, respectively, as marked), with various stellar mass cuts (above $10^{7}$ M in solid lines and $10^{9}$ M dashed). Systems with no dynamical friction are shown at $\mathrm{ln}(\Lambda)$ = 0. The lightened lines show the position of the median splashback radius from the default variable implementation of dynamical friction (as used in Figures 2 to 11.) The shaded bands show the 16-84 percentile range for all data.
Appendix C Subhalo angular momentum
Although a thorough examination of angular momentum in clusters is beyond the scope of our study, here we briefly discuss how dynamical friction affects the angular momentum of galaxies in our sample.
Figure 13 shows the ratio of angular momentum $L\equiv M_{*}rv_{\rm tan}$ for a matched sample of galaxies with and without dynamical friction. The sample consists of all galaxies that have the same residence time in both simulations. We expect to lose some galaxies due to variations in the infall time between the two simulations. In particular, this method is likely to exclude galaxies where the residence time is significantly altered by dynamical friction, potentially underestimating the impact of DF, however by comparing a matched sample, we are able to target dynamical friction directly.
The impact of dynamical friction on galaxy orbits may vary widely. For galaxies which pass very close to the cluster center, dynamical friction may radically affect their orbits and angular momentum. Figure 1 shows a few extreme cases of these orbital changes. Most galaxies will experience a more moderated effect. Fig. 13 shows the population level impact of DF on angular momentum.
Figure 13: Median ratio of angular momentum of subhaloes after 10 Gyr evolved with and without dynamical friction, as a function of their residence time in the cluster. Here, we show only the least massive cluster bin, and separate galaxies by stellar mass. Note that we impose a maximum stellar mass cut as well as minimum, unlike the rest of the analysis.
As expected, more massive galaxies lose more angular momentum when dynamical friction is enabled. This decrease in $L$ and associated energy loss is also a function of residence time; galaxies that have resided in the cluster for longer have lost more angular momentum to dynamical friction. At its maximum, this effect reaches ${\sim}10\%$ for the most massive galaxies in the lowest mass clusters, consistent with our findings for the splashback radius in Fig. 6. In the highest mass clusters ($M_{\rm 200,mean}>10^{14.5}\rm M_{\odot}$) the median galaxy does not lose more than 0.5% of its angular momentum from dynamical friction.